# Cycle Consistency as Reward: Learning Image-Text Alignment without Human Preferences

### v1

@@ Abstract
Learning alignment between language and vision is a fundamental challenge, especially as multimodal data becomes increasingly detailed and complex. Existing methods often rely on collecting human or AI preferences, which can be costly and time-intensive. We propose an alternative approach that leverages cycle consistency as a supervisory signal. Given an image and generated text, we map the text back to image space using a text-to-image model and compute the similarity between the original image and its reconstruction. Analogously, for text-to-image generation, we measure the textual similarity between an input caption and its reconstruction through the cycle. We use the cycle consistency score to rank candidates and construct a preference dataset of 866K comparison pairs. The reward model trained on our dataset outperforms state-of-the-art alignment metrics on detailed captioning, with superior inference-time scalability when used as a verifier for best-of-$N$ sampling. Furthermore, performing DPO and Diffusion DPO using our dataset enhances performance across a wide range of vision-language tasks and text-to-image generation. Our dataset, model, and code are publicly released at https://cyclereward.github.io/.
Figure 1: Method overview. (a) Given an input image $x$, we generate multiple candidate captions $F_{i}(x)$, $F_{j}(x)$ using different captioning models. Each caption is mapped back to the image domain via a text-to-image model $G$, and compared against the original image. Captions whose reconstructions $G(F(x))$ are more similar to the original image are preferred; those with low similarity are rejected. (b) These comparison pairs are used to train a reward model, which learns to assign higher scores to preferred captions. We apply the same process for text-to-image generation.
@@ 1 Introduction
Aligning language and vision remains a central challenge in multimodal learning. As input data becomes increasingly complex and detailed, generative models often produce misaligned outputs — vision-language models often hallucinate descriptions not grounded in the image [50], while diffusion models can generate images that misrepresent quantities, attributes, or object relationships in text prompts [33, 40]. Existing alignment metrics primarily focus on short captions or prompts [91, 40, 89], limiting their effectiveness for evaluating dense alignment. Common strategies for improving alignment include reinforcement learning from human feedback (RLHF) [65] or direct preference optimization (DPO) [70], which requires collecting high-quality human preferences [91, 40, 89, 88] or AI feedback [48] from proprietary models (e.g., GPT-4V [62]), which is prohibitively expensive.
Comparing images and text is inherently challenging, even more so with detailed text. However, the comparison becomes much easier when we map text back into image space (Figure 2). As the generated text becomes more descriptive and accurate, the reconstructed image better resembles the original image. This idea of cycle consistency [39, 81, 106] has been used as a metric to evaluate image-to-text generation [32, 23] and to optmize diffusion models [4]. However, these approaches compute cycle consistency on-the-fly using large pre-trained models, which is prohibitively slow and often not differentiable.
In this paper, we introduce a reward model trained on preferences derived from cycle consistency. Given an image-to-text mapping $F:X\rightarrow Y$ and a backward text-to-image mapping $G:Y\rightarrow X$, we define cycle consistency score as the similarity between the original input $x$ and its reconstruction $G(F(x))$. We use the cycle consistency score as a proxy for preferences, where a higher score indicates a preferred output. This provides a more scalable and cheaper signal for learning image-text alignment compared to human supervision.
We construct CyclePrefDB, a large-scale preference dataset comprising 866K comparison pairs from 11 image-to-text models and 4 text-to-image models. CyclePrefDB provides significantly denser text than typical text-to-image datasets (Table 1), while remaining within the 77-token limit of text-to-image models to compute cycle consistency. We train a reward model on this dataset, called CycleReward, which proves effective both as an alignment metric and for best-of-$N$ optimization. It achieves state-of-the-art performance for detailed captioning and competitive performance in text-to-image synthesis. Finally, applying direct preference optimization (DPO) [70, 85] using our dataset enhances a wide range of vision-language and text-to-image generation tasks without requiring any human supervision.
@@
  • CyclePrefDB, a cycle-consistent preference dataset of 866K comparison pairs spanning image-to-text and text-to-image generation, specifically designed for dense captions and prompts.
  • CycleReward, a reward model trained on our dataset, which is effective both as an alignment metric and a verifier for best-of-$N$ optimization.
  • Our ablation study finds that image-to-text decoders with stronger language models and text-to-image decoders lead to better alignment. Additionally, using similarity metrics that model human perception improve alignment.
  • Demonstration of DPO using our CyclePrefDB dataset. This leads to improvements on a wide range of vision-language tasks and text-to-image generation tasks.
Figure 2: What do cycle consistency preferences look like? We visualize examples from our dataset, where cycle consistency determines preferences over comparison pairs. Preferred samples are in blue and rejected samples are in red. Image-to-text generation (left): The preferred caption provides a fine-grained, detailed description resulting in a faithful reconstruction of the original image, whereas the rejected caption is short and vague, producing a reconstruction far from the original image. Text-to-image generation (right): Images that capture fine-grained details of the input prompt produce better text reconstructions, resulting in higher cycle consistency. See Appendix C for more examples.
@@ 2 Related Work
Image-text alignment. Image-text alignment metrics can be classified in two ways - as reference-based, which require comparison with ground truth text, or as reference-free, which compute alignment based solely on the provided image and text. Reference-based metrics include BLEU [66], CIDEr [84], and METEOR [41] which measure linguistic similarity between candidate and reference captions. These methods do not generalize well to texts which vary in style and syntax from the reference caption. Recent approaches such as SPICE [1], CAPTURE [16], and DCScore [92] decompose the candidate text into scene graphs or basic information units which are then compared to ground truth labels. Although recent reference-based metrics are more flexible and thorough, they are limited by lack of differentiability, slow runtime, and, most importantly, they require a reference, which means they are not suitable as an objective function.
Reference-free metrics come in a variety of forms. Many approaches adapt pre-trained CLIP [69] image and text encodings [29, 18, 37, 98] while others collect human preferences to train reward models [40, 91, 89, 88]. Some recent methods query a large pre-trained model to directly evaluate alignment  [52, 7, 42, 90]. Although current metrics increasingly align with human preferences for visio-linguistic reasoning and text-to-image evaluation, many of these methods fail to evaluate longer, more descriptive captions effectively. Most similar to our method, Image2Text2Image [32] computes image captioning performance by leveraging text-to-image generation to produce reconstructed images given text captions. The final score is the reconstruction error between the original and generated image’s DINOv2 [64] or CLIP [69] features. DDPO [4] also includes a similar text-to-image-to text reconstruction score to optimize diffusion models. These pipelines match our dataset collection process outlined in Section 3. However, our method uses cycle consistency scores to train a reward model with the benefit of inference speed, differentiability for downstream applications, and better performance.
Detailed captioning. Image-to-text models can produce comprehensive descriptions [24, 61] by scaling the language model [53, 54] and training on semantically rich synthetic captions [46, 47, 78, 53, 54]. Despite these growing model capabilities, little attention has been given to evaluating descriptive captions. DetailCaps-4870 [16] and DeCapBench [92] address this issue. DetailCaps-4870 [16] evaluates metrics on their ability to measure alignment for detailed descriptions, whereas DeCapBench evaluates image-to-text model results on detailed captioning using their reference-based metric DCSCORE. Our reward model provides a fast, differentiable, and reference-free approach to measuring alignment for descriptive texts.
@@
There are several techniques to align model outputs with human preferences [79, 65] at training [75, 70, 77] or test time [79, 60, 38]. These approaches have been applied mostly to large language models and recently in vision-language models [96, 80, 97] and diffusion models [85, 4, 68]. Text-to-image alignment metrics such as Human Preference Score (HPS) [89, 88], PickScore [40], and ImageReward [91] all collect human preferences to train a reward model. Other methods such as VLFeedback [48] substitute human feedback by using foundation models (e.g., GPT-4V) to annotate preferences [92, 48, 102, 97] and perform Direct Preference Optimization (DPO) [70] on their dataset. Our method collects preferences from a new signal: cycle consistency, which is cheaper and more easily scalable. We apply our dataset both to reward modeling and preference learning via DPO, exhibiting competitive performance with models trained on human labels.
@@ 3.1 Cycle Consistency as Preferences
Our goal is to learn preferences for image-text alignment without relying on human annotations. Prior approaches often use human [40, 89, 91] or GPT-4V [48] preferences to rank the quality of generated captions or images. Instead, we propose to derive preferences from cycle consistency. Given an image-to-text mapping $F:X\rightarrow Y$, we measure how well a text $F(x)$ aligns with an image $x$ by measuring how well it can reconstruct $x$ through a backward mapping $G:Y\rightarrow X$. We define cycle consistency score for $F(x)$ conditioned on $x$ as:
@@
Similarly, for a text-to-image mapping $G:Y\rightarrow X$, we measure how well an image $G(y)$ aligns with text $y$ by using a backward mapping $F:X\rightarrow Y$. We define the cycle consistency score for $G(y)$ conditioned on $y$ as:
@@
While prior work [23, 32] uses this score directly as an alignment metric, we learn preferences from a large pool of comparisons. Given triplets $(x,y_{i},y_{j})$ and $(y,x_{i},x_{j})$, we convert the cycle consistency scores into pairwise preferences:
@@
Table 1: Key differences of preference datasets. Existing preference datasets often use human or GPT-4V annotations for supervision, whereas we leverage cycle consistency for preference annotation. We provide comparison pairs for both image-to-text (I2T) and text-to-image (T2I) generation tasks. CyclePrefDB features significantly denser text than typical T2I datasets, while remaining within token limits (77 tokens) of text-to-image models. VL denotes vision-language tasks.
@@ 3.2 Dataset Generation
We design our dataset to capture dense alignment between images and text, focusing on captioning images with rich descriptions and generating images from longer, detailed prompts. To this end, we use the train split of Densely Captioned Images (DCI) dataset [83] for input images and texts. It contains 7.6K image-text pairs featuring high-resolution images annotated with dense, detailed captions. Due to prompt length constraints of text-to-image models, we use sDCI, a summarized version of DCI to fit within 77 tokens. For details and visualizations see Appendix C.
Image-to-text generation. Given image $x$, we first obtain multiple candidate text descriptions $\{y_{1},...,y_{n}\}$ of varying quality. In practice, we use 11 image-to-text models trained on different datasets and scales: BLIP2 (T5-XXL) [47], LLaVA-1.5 (7B, 13B) [55], LLaVA-1.6 (7B, 34B) [54], LLaVA-OneVision (0.5B, 7B) [44], and InternVL2 (2B, 8B, 26B, 40B) [63, 9]. As reward modeling is inherently contrastive, we deliberately include older models that produce short, hallucinated captions as negative examples alongside newer models to maximize text diversity. We specifically instruct the models to generate rich, descriptive captions, using the prompt recommended by the model distributor (Appendix C). We use greedy sampling with a maximum token length of 77, i.e., maximum prompt length supported by the text-to-image models. We fix the backward mapping $F$ as LLaVA-1.5-13B to compute $s(x\rightarrow y)$.
Text-to-image generation. Given a text prompt $y$, we generate a set of image candidates $\{x_{1},...,x_{n}\}$, similarly using 4 text-to-image models: Stable Diffusion 1.5 [72], Stable Diffusion XL [67], Stable Diffusion 3 [20], and FLUX (Timestep-distilled) [5]. Similarly, we select models with varying prompt-following capabilities to maximize diversity of generated images. We use three random seeds to generate the images, creating 12 candidate images per prompt. We fix the backward mapping $G$ as Stable Diffusion 3 to compute $s(x\rightarrow y)$.
@@ 3.3 Reward Modeling
The generality of cycle-consistent preferences allows us to train a reward model in multiple ways. We explore three variants: (1) CycleReward-I2T: trained with image-to-text preferences $s(x\rightarrow y)$, (2) CycleReward-I2T: trained with text-to-image preferences $s(y\rightarrow x)$, and (3) CycleReward-Combo: jointly trained with both datasets.
@@
Given a dataset of image-to-text comparison pairs $(x,y_{i},y_{j})$, where image $x$ is paired with a preferred text $y_{i}$ and rejected text $y_{j}$, the loss [65, 79] is formulated as:
@@
where $r_{\theta}(x,y)$ is the scalar output of the reward model.
Similarly, given a dataset of text-to-image comparison pairs $(y,x_{i},x_{j})$, where text $y$ is paired with a preferred image $x_{i}$ and rejected image $x_{j}$, the loss is formulated as:

### v2

@@ Abstract
Measuring alignment between language and vision is a fundamental challenge, especially as multimodal data becomes increasingly detailed and complex. Existing methods often rely on collecting human or AI preferences, which can be costly and time-intensive. We propose an alternative approach that leverages cycle consistency as a supervisory signal. Given an image and generated text, we map the text back to image space using a text-to-image model and compute the similarity between the original image and its reconstruction. Analogously, for text-to-image generation, we measure the textual similarity between an input caption and its reconstruction through the cycle. We use the cycle consistency score to rank candidates and construct a preference dataset of 866K comparison pairs. The reward model trained on our dataset, CycleReward, outperforms state-of-the-art alignment metrics on detailed captioning, with superior inference-time scalability when used as a verifier for Best-of-$N$ sampling, while maintaining speed and differentiability. Furthermore, performing DPO and Diffusion DPO using our dataset enhances performance across a wide range of vision-language tasks and text-to-image generation. Our dataset, model, and code are publicly released at https://cyclereward.github.io/.
Figure 1: Method overview. (1) Given an input image $x$, we generate multiple candidate captions $F_{i}(x)$, $F_{j}(x)$ using different captioning models. Each caption is mapped back to the image domain via a text-to-image model $G$, and compared against the original image. Captions whose reconstructions $G(F(x))$ are more similar to the original image are preferred; those with low similarity are rejected. (2) These comparison pairs are used to train a reward model, which learns to assign higher scores to preferred captions. We apply the same process for text-to-image generation. (3) At test time, the trained reward model outputs alignment scores for arbitrary image-text pairs.
@@ 1 Introduction
Measuring image–text alignment is a central problem in multimodal learning, where the goal is to learn a metric $d(x,y)$ that quantifies the correspondence between an image $x$ and text $y$. Such metrics are essential for evaluating vision–language and text-to-image models [29, 51, 91, 40, 89] and improving model alignment through test-time optimization [10, 79, 57] or reinforcement learning from human feedback (RLHF) [64]. However, existing metrics typically rely on high-quality human preference data [91, 40, 89, 88], which are expensive to collect and difficult to scale. Moreover, most of these datasets focus on short text [91, 40, 89], limiting their ability to assess alignment for longer and more complex text. Another method uses AI feedback [48] from proprietary models (e.g., GPT-4V [61]), which are costly, closed-source, and rate-limited via APIs, limiting long-term accessibility and scalability.
Comparing images and text is inherently challenging, especially with longer, detailed text. However, the comparison becomes much easier when we map text back into image space. As shown in Figure 1, more descriptive and accurate texts lead to reconstructed images that better resemble the original images. This idea of cycle consistency [39, 81, 106] has been used as a metric to evaluate image-to-text generation [32, 23] and optimize diffusion models [4]. However, these approaches compute cycle consistency on-the-fly using large pre-trained models, which is prohibitively slow and often not differentiable.
We introduce CycleReward, a reward model trained on preferences derived from cycle consistency. Given an image-to-text mapping $F:X\rightarrow Y$ and a backward text-to-image mapping $G:Y\rightarrow X$, we define cycle consistency score as the similarity between the original input $x$ and its reconstruction $G(F(x))$. In the opposite direction, we can compare reconstructed text $F(G(y))$ with input text $y$. We use the cycle consistency score as a proxy for preferences, where a higher score indicates a preferred output. This provides a more scalable and cheaper signal for learning alignment compared to human supervision. We create a large-scale preference dataset, CyclePrefDB, comprising 866K comparison pairs from 11 image-to-text models and 4 text-to-image models. It contains significantly denser text than typical text-to-image datasets (Table 1), while fitting within the 77-token limit of text-to-image models. Trained on this dataset, CycleReward is a fast, differentiable metric for image-text alignment, particularly for longer text.
We evaluate CycleReward’s ability to evaluate and enhance image-text alignment across two tasks: detailed captioning and text-to-image generation. We find that it is effective both metric for evaluation and Best-of-$N$ optimization. It achieves state-of-the-art performance for detailed captioning and performs competitively text-to-image synthesis. Finally, applying direct preference optimization (DPO) [69, 85] using CyclePrefDB enhances a wide range of vision-language and text-to-image generation tasks without requiring any human supervision.
@@
  • CyclePrefDB, a cycle consistency based preference dataset of 866K comparisons for image-to-text and text-to-image generation, specifically for longer texts.
  • CycleReward, a reward model trained on our dataset, which is effective as a fast, differentiable alignment metric and a verifier for Best-of-$N$ sampling for longer text.
  • Our ablation study finds that image-to-text decoders with stronger language models lead to better alignment. Additionally, using similarity metrics that model human perception improves alignment.
  • Demonstration of DPO using our CyclePrefDB dataset, leading to improvements on a wide range of vision-language and text-to-image generation tasks.
Figure 2: What do cycle consistency preferences look like? We visualize comparison pairs from our dataset, where cycle consistency determines preferences. Preferred samples are in blue and rejected samples are in red. Image-to-text generation (left): The preferred caption provides a fine-grained description resulting in a faithful reconstruction of the original image, whereas the rejected caption is short and vague, producing a reconstruction far from the original image. Text-to-image generation (right): Images that capture fine-grained details of the input prompt produce better text reconstructions, resulting in higher cycle consistency. See Appendix C for more examples.
@@ 2 Related Work
Image-text alignment. Image-text alignment metrics can be classified either as reference-based, which require comparison with ground truth text, or as reference-free, which compute alignment based solely on the provided image and text. Reference-based metrics include BLEU [65], CIDEr [84], and METEOR [41] which measure linguistic similarity between candidate and reference captions, but often do not generalize well to texts which vary in style and syntax from the reference caption. Recent approaches such as SPICE [1], CAPTURE [16], and DCScore [92] decompose the candidate text into scene graphs or basic information units which are then compared to ground truth labels. Although these recent metrics are more flexible and thorough, they are limited by lack of differentiability, slow runtime, and, most importantly, they require a reference, which means they are not suitable as an objective function.
Reference-free metrics come in a variety of forms. Many approaches adapt pre-trained CLIP [68] image and text encodings [29, 18, 37, 98] while others collect human preferences to train reward models [40, 91, 89, 88]. Some recent methods query a large pre-trained model to directly evaluate alignment  [51, 7, 42, 90, 70]. Although current metrics increasingly align with human preferences for visio-linguistic reasoning and text-to-image evaluation, many of these methods fail to evaluate longer, more descriptive captions effectively. Most similar to our method, Image2Text2Image [32] computes image captioning performance by leveraging text-to-image generation to produce reconstructed images given text captions. The final score is the reconstruction error between the original and generated image’s DINOv2 [63] or CLIP [68] features. DDPO [4] also includes a similar text-to-image-to text reconstruction score to optimize diffusion models. These pipelines match our dataset collection process outlined in Section 3. However, our method uses cycle consistency scores to train a reward model with the benefit of inference speed, differentiability for downstream applications, and better performance.
Detailed captioning. Image-to-text models can produce comprehensive descriptions [24, 60] by scaling the language model [52, 53] and training on semantically rich synthetic captions [46, 47, 78, 52, 53]. Despite growing model capabilities, little attention has been given to evaluating descriptive captions. Addressing this issue, DetailCaps-4870 [16] evaluates image-text alignment metrics on detailed descriptions, whereas DeCapBench [92] evaluates image-to-text models on detailed captioning using their reference-based metric DCSCORE. Our reward model provides a fast, differentiable, and reference-free approach to measuring alignment for descriptive texts.
@@
There are many techniques to align model outputs with human preferences [79, 64] at training [75, 69, 77] or test time [79, 59, 38]. These approaches have been applied mostly to large language models and recently to vision-language models [96, 80, 97] and diffusion models [85, 4, 67]. Text-to-image alignment metrics such as Human Preference Score (HPS) [89, 88], PickScore [40], and ImageReward [91] all collect human preferences to train a reward model. VLFeedback [48] substitutes human feedback by using foundation models (e.g., GPT-4V) to annotate preferences [92, 48, 102, 97] and applies Direct Preference Optimization (DPO) [69] with their dataset. Our method collects preferences from a new signal: cycle consistency, which is cheaper and more easily scalable. We apply our dataset both to reward modeling and preference learning via DPO, exhibiting competitive performance with models trained on human labels.
@@ 3.1 Cycle Consistency as Preferences
Our goal is to learn preferences for image-text alignment without relying on human annotations. Prior approaches often use humans [40, 89, 91] or GPT-4V [48] to rank the quality of generated captions or images. Instead, we propose to derive preferences from cycle consistency. Given image-to-text mapping $F:X\rightarrow Y$, we measure how well text $F(x)$ aligns with image $x$ by measuring how well backward mapping $G:Y\rightarrow X$ can reconstruct $x$. We define cycle consistency score for $F(x)$ conditioned on $x$ as:
@@
Similarly, for text-to-image mapping $G:Y\rightarrow X$, we measure how well image $G(y)$ aligns with text $y$ by using a backward mapping $F:X\rightarrow Y$. We define the cycle consistency score for $G(y)$ conditioned on $y$ as:
@@
While prior work [23, 32] uses this score directly as an alignment metric, we learn alignment from a large pool of comparisons. Given triplets $(x,y_{i},y_{j})$ and $(y,x_{i},x_{j})$, we convert cycle consistency scores into pairwise preferences:
@@
Table 1: Key differences of preference datasets. Existing preference datasets use human or GPT-4V annotations for supervision, whereas we label preferences with cycle consistency. We provide comparison pairs for both image-to-text (I2T) and text-to-image (T2I) tasks. CyclePrefDB features significantly denser text than typical T2I datasets, while remaining within token limits (77 tokens) of text-to-image models. VL denotes vision-language tasks.
@@ 3.2 Dataset Generation
We design our dataset to capture alignment between images and dense text, focusing on captioning images with rich descriptions and generating images from longer, detailed prompts. To this end, we use the train split of Densely Captioned Images (DCI) dataset [83] for input images and texts. It contains 7.6K image-text pairs featuring high-resolution images annotated with dense captions. Due to prompt length constraints of text-to-image models, we use sDCI, a summarized version of DCI to fit within 77 tokens. See Appendix C for details and visualizations.
Image-to-text generation. Given image $x$, we first obtain multiple candidate text descriptions $\{y_{1},...,y_{n}\}$ of varying quality. In practice, we use 11 image-to-text models trained on different datasets and scales: BLIP2 (T5-XXL) [47], LLaVA-1.5 (7B, 13B) [54], LLaVA-1.6 (7B, 34B) [53], LLaVA-OneVision (0.5B, 7B) [44], and InternVL2 (2B, 8B, 26B, 40B) [62, 9]. As reward modeling is inherently contrastive, we deliberately include older models that produce short, hallucinated captions as negative examples alongside newer models to maximize text diversity. We specifically instruct the models to generate rich, descriptive captions, using the prompt recommended by the model distributor (Appendix C). We use greedy sampling with a maximum token length of 77, i.e., maximum prompt length supported by the text-to-image models. We fix the backward mapping $G$ as Stable Diffusion 3 to compute $s(x\rightarrow y)$.
Text-to-image generation. Given a text prompt $y$, we generate a set of image candidates $\{x_{1},...,x_{n}\}$ using 4 text-to-image models: Stable Diffusion 1.5 [72], Stable Diffusion XL [66], Stable Diffusion 3 [20], and FLUX (Timestep-distilled) [5]. Similarly, we select models with varying performance to maximize diversity of generated images. We use three random seeds to generate the images, creating 12 candidate images per prompt. We fix the backward mapping $F$ as LLaVA-1.5-13B to compute $s(y\rightarrow x)$.
@@ 3.3 Reward Modeling
The generality of cycle-consistent preferences allows us to train a reward model in multiple ways. We explore three variants: (1) CycleReward-I2T: trained with image-to-text preferences $s(x\rightarrow y)$, (2) CycleReward-I2T: trained with text-to-image preferences $s(y\rightarrow x)$, and (3) CycleReward-Combo: jointly trained on both datasets.
@@
Given a dataset of image-to-text comparisons $(x,y_{i},y_{j})$, where image $x$ is paired with preferred text $y_{i}$ and rejected text $y_{j}$, the loss is formulated as:
@@
where $r_{\theta}(x,y)$ is the scalar output of the reward model [64, 79]. Similarly, given a dataset of text-to-image comparison pairs $(y,x_{i},x_{j})$, where text $y$ is paired with a preferred image $x_{i}$ and rejected image $x_{j}$, the loss is formulated as:
@@
4 Cycle Consistency and Human Preferences
Does cycle consistency align with human preferences? We measure the agreement rate between cycle consistency and human preferences on detailed captioning and text-to-image generation. For detailed captioning, we compare to human preferences from RLHF-V [96] and POVID [105] datasets. For text-to-image generation, we compare to HPDv2 [88], Pick-a-Pic v2 [40], and ImageRewardDB [91]. For each dataset, we sample 1K random binary comparison pairs. We compare human labels to raw cycle consistency scores, $s(x\rightarrow y)$ and $s(y\rightarrow x)$, as well as our trained reward models. We also compare against GPT-4o [60] annotations, as they have been shown effective for preference learning [48].
Table 2 shows that CycleReward achieves the highest agreement with human annotations, with CycleReward-Combo having the highest average agreement rate of 65%. While GPT-4o annotations on detailed captioning align more closely with humans, agreement drops significantly on text-to-image generation, with as low as 24.84% on ImageRewardDB. In contrast, raw cycle consistency has a consistent agreement rate across both tasks. Training a reward model with cycle consistency further improves alignment, demonstrating the effectiveness of distilling cycle-consistent preferences into a learned reward model.
While we compare against human preferences, our aim is not to mimic them. Instead, we aim to learn image-text alignment, and demonstrate that cycle consistency is an effective proxy—achieving strong results without collecting any human labels, as shown in the following sections.